# Thomassen’s proof and Filippov’s proof of the Weak Jordan Theorem

### 05 Apr 2026 v1

@@ 3 Filippov’s proof of the WJT
Let $c\in S(f)$. As in Step 1 of Thomassen’s proof, we denote by $\ell(c)$ the half-line going up from $c$ and consider the geometric intersection $P(c)$ of $\ell(c)$ and $f[I]$. For $P(c)=\emptyset$ we set $N(c)=0$. Let $P(c)\neq\emptyset$. The complication, but it is not really a complication, is that now $f$ need not be injective. Instead of $\ell(c)\cap f[I]$ we work with the inverse image of this set. Let $a=a_{0}<a_{1}<\dots<a_{k}=b$ be the partition of the domain $[a,b]$ of $f$. For every $i=1,2,\dots,k$, the segment $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ is either disjoint to $\ell(c)$ or $s_{i}\cap\ell(c)=\{p\}$ for a single point $p$ or $s_{i}\subset\ell(c)$. It follows that
@@
is a finite disjoint union of components $z$ such that every $z$ is either a subinterval $a_{i},a_{j}]$ ($\subset[a,b]$) with $0\leq i<j\leq k$, where for non-closed $f$ the first and last inequality is strict, or an isolated point in $(a,b)$. We call zzsimple component if the values $f(t)$ for $t$ immediately before and after $z$ lie on different sides of $\ell(c)$. Else, if they lie on the same side of $\ell(c)$, we call zzdouble component. We define
@@
We prove that the parity map $N$ is continuous on $S(f)$. Let $c\in S(f)$. We take a $\delta>0$ small enough such that the ball $B=B(c,\delta)\subset S(f)$ and that no vertical line that intersects $B$ contains a corner of $f[I]$ distinct from those appearing in $P(c)$. Let $c^{\prime}\in B$ with $c^{\prime}\neq c$. As we move $c$ to $c^{\prime}$, every simple component of $P(c)$ is preserved and every double component is either fixed or disappears or decays in two simple (point) components. It follows that $N(c)=N(c^{\prime})$. So $N$ is continuous at $c$ and on $S(f)$.
@@
Step 2. Two properties of the parity map $N$. The first property is the additivity of $N(c,f)$ in the variable $f$. Let $f_{1},f_{2}\colon I=[a,b]\to\mathbb{R}^{2}$ be two non-closed PL maps such that $f_{1}(a)=f_{2}(b)$ and $f_{1}(b)=f_{2}(a)$, so that the PL map $f_{3}=f_{1}+f_{2}$ is closed. Let $c\in\mathbb{R}^{2}\setminus(f_{1}[I]\cup f_{2}[I])$ be such that $f_{1}(a)_{x}<c_{x}<f_{1}(b)_{x}$. Then
@@
The proof is short, but it takes some effort to spell it out rigorously: if $S_{i}$ is the set of simple components in $P(c,f_{i})$, $i=1,2,3$, then there is a bijection between $S_{3}$ and $(S_{1}\times\{0\})\cup(S_{2}\times\{1\})$ (we leave the details of the bijection to the reader as an exercise).
@@
To prove it, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $I$ and the components $z_{1}<z_{2}<\dots<z_{l}$ of
@@
As we know, every component $z_{i}$ is an interval $a_{r_{i}},a_{s_{i}}]$ ($\subset[a,b]$) with $0<r_{i}<s_{i}<k$ or an isolated point $\{t_{i}\}$ with $t_{i}\in(a,b)$. It is clear that $P(c,f)\neq\emptyset$ and $l\geq 1$ (Proposition 1.18). We call the $l+1$ open subintervals
@@
of $(a,b)$ gaps. It follows from the assumption on $c$ that for every gap $g_{i}$ and $t\in g_{i}$, all values $f(t)$ lie on the left side ($L$) of $\ell(c)$, or on the right side ($R$). Accordingly, we associate with $P(c,f)$ a word $w_{0}w_{1}\dots w_{l}\in\{L,R\}^{l+1}$. For $i=1,2,\dots,l$, the transitions $w_{i-1}w_{i}$ with $w_{i-1}=w_{i}$ correspond to the double components of $P(c,f)$, and those with $w_{i-1}\neq w_{i}$ to the simple components. Since $w_{0}\neq w_{l}$, the number of the latter transitions is odd.
Again, the reason for $N(c,f)=1$ is quite intuitive — any PL map going from one side of a line $\ell$ to the other side has to cross $\ell$ an odd number of times — but if we want to go beyond the intuitive level to a more rigorous level, which is what we want to do here, it takes some effort.
@@
Step 3. Some two points in the complement of any circuit are separated. Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $D=\mathbb{R}^{2}\setminus f[I]$. We may assume that $f[I]$ lies to the right of the $y$-axis. Let $l\in f[I]$ and $p\in f[I]$ be a leftmost and a rightmost point of $f[I]$, respectively. These points exist by Propositions 1.4 and 1.5, and they divide $f$ in two arcs $f_{1},f_{2}\colon I\to\mathbb{R}^{2}$ such that $f_{1}^{0}\cap f_{2}^{0}=\emptyset$, $f_{1}[I]\cup f_{2}[I]=f[I]$ and $\mathrm{ep}(f_{1})=\mathrm{ep}(f_{2})=\{l,p\}$ (Proposition 1.17). Clearly, $l_{x}<p_{x}$ (Proposition 1.16). Let $\ell$ be any vertical line with $x$-coordinate in $(l_{x},p_{x})$. Let $a\in f[I]\cap\ell$ be the highest intersection point of the circuit and the line (Propositions 1.4 and 1.5). We may assume that $a\in f_{1}^{0}$. Similarly, let $b\in f_{1}^{0}\cap\ell$ be the lowest intersection point of the arc $f_{1}$ and the line. Let $c\in\ell$ lie below $b$ but let
$0<d_{2}(c,\,b)<d_{2}(b,f_{2}[I])\,.$
@@
Then $c\in D$; the point $c$ exists by Proposition 1.8, by which $d_{2}(b,f_{2}[I])>0$. We show that for every PL map $g\colon I\to\mathbb{R}^{2}$ that joins $c\in D$ to the origin $\overline{0}=\langle 0,0\rangle\in D$ ($f[I]$ lies to the right of the $y$-axis), $g[I]\cap f[I]\neq\emptyset$. In view of Propositions 1.6 and 1.15 it means that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is disconnected.
@@
Suppose for the contrary that $g\colon I\to\mathbb{R}^{2}$ is a PL map joining $c$ to $\overline{0}$ such that $g[I]\cap f[I]=\emptyset$. Let $\kappa$ be the vertical half-line going down from $c$ and let
$\lambda=c\,b\cup f_{3}[I]\cup\ell(a)\,$,
@@
where $cb$ is the straight segment joining these two points, $f_{3}$ is a subarc of $f_{1}$ joining $b$ to $a$ (we omit $f_{3}[I]$ if $b=a$) and, as before, $\ell(a)$ is the half-line going up from $a$. Let $h>0$ be the minimum of the three positive distances
$d_{2}(g[I],\,f[I])>0,\ d_{2}(f_{1}[I],\,\kappa)>0\,\text{ and }\,d_{2}(f_{2}[I],\,\lambda)>0\,.$
@@
The first distance is positive by the assumption that $g[I]\cap f[I]=\emptyset$ and by Proposition 1.8. The latter two distances are positive by Proposition 1.9.
@@
We approximate the given circuit $f\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ by a sufficiently fine closed PL map. By Proposition 1.10, there exists a partition $\alpha=a_{0}<a_{1}<\dots<a_{k}=\beta$ of $\alpha,\beta]$ such that $\{a,b,l,p\}\subset\{f(a_{i})\colon\;i=0,1,\dots,k\}$ and
@@
Joining the consecutive points $f(a_{i-1})$ and $f(a_{i})$ for $i=1,2,\dots,k$ ($f(a_{0})=f(a_{k})$) by straight segments, we obtain a closed PL map $f_{4}\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ such that all corners of $f_{4}[I]$ lie on $f[I]$ and include the four points $a$, $b$, $l$, and $p$. Also, for every $z\in f_{4}[I]$ we have $d_{2}(z,f[I])<\frac{1}{2}h$. Let $f_{5}$, $f_{6}$, and $f_{7}$ be the non-closed PL sub-maps of $f_{4}$ spanned by the corners of $f_{4}[I]$ lying in $f_{1}[I]$, $f_{2}[I]$, and $f_{3}[I]$, respectively. The above definition of $h$ and the $\frac{1}{2}h$-approximation of $f_{4}[I]$ by $f[I]$ imply that
@@
The first equality follows from the fact that since $f[I]$ lies to the right of the $y$-axis, so does $f_{4}[I]$. The second equality follows from Step 1 because $f_{4}[I]\cap g[I]=\emptyset$. The third equality follows from the first property of $N$ in Step 2 because (we can define $f_{5}$ and $f_{6}$ such that) $f_{4}=f_{5}+f_{6}$. We have $N(c,f_{5})=1$ by the second property of $N$ in Step 2. Finally, $N(c,f_{6})=N(a,f_{6})$ by Step 1 because $(cb\cup f_{7}[I])\cap f_{6}[I]=\emptyset$, and $N(a,f_{6})=0$ by the definition of $N$ because $\ell(a)\cap f_{6}[I]=\emptyset$.
@@ 4 Concluding remarks
Jordan’s proof of his theorem in [7] was questioned in the past, and it was claimed that the first correct proof was given by Veblen [10]. Hales reversed these assessments in [5] where he showed that Jordan’s proof is basically correct, when some gaps are filled, but that Veblen’s proof has serious problems. Fillipov’s article [3] is preceded by the article [11] due to A. I. Vol’pert, which contains a similar proof of the JT. The footnote on the first page of [3] reads “After this work was completed, the author learned that a proof of Jordan’s theorem, very close to the one in this article, was somewhat earlier found by A. I. Vol’pert (Lviv)”. Authors of these two proofs of JT, Aleksei F. Filippov (1923–2006) and Aizik I. Volpert (Vol’pert) (1923–2006), were excellent mathematicians; see [2] and [1].
What are, or should be, the main results of a mathematician’s work today? I think: rigorous proofs — that is, routinely, algorithmically, formalizable proofs — of mathematical results, of theorems; see, for example, Hamami [6]. It is sometimes claimed that explanations, generalizations, abstractions, analogies, connections, motivations, applications, and who knows what else are more important than proofs. In my opinion, this is incorrect, especially today. The mentioned things are indeed important, but if they are not supported by strong and correct theorems, they are just castles in the air. The onset of formalized mathematics ([4] describes formalization of Thomassen’s proof of the JT) and AI changed everything, compared to how the relation of informal and formal proofs was viewed and thought about, say in the previous millennium. In this article, we tried to lift Thomassen’s proof and Filippov’s proof on this modern, more demanding level of rigour.
I lectured on Thomassen’s proof of the WJT in the course Mathematical Analysis 3 in the years 2023–2026. For the next year I plan to switch to the proofs of the JT in [3, 11].

### 19 Apr 2026 v3

@@ 3 Filippov’s proof of the WJT
Let $c\in S(f)$. As in Claim 1 of Thomassen’s proof, we denote by $\ell(c)$ the half-line going up from $c$ and consider the geometric intersection $P(c)$ of $\ell(c)$ and $f[I]$. For $P(c)=\emptyset$ we set $N(c)=0$. Let $P(c)\neq\emptyset$. Since now $f$ need not be injective, we consider instead of $\ell(c)\cap f[I]$ the preimage of this set. Let $a=a_{0}<a_{1}<\dots<a_{k}=b$ be the partition of the domain $[a,b]$ of $f$. For every $i=1,2,\dots,k$, the segment $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ of $f[I]$ is either disjoint to $\ell(c)$ or $s_{i}\cap\ell(c)=\{p\}$ for a single point $p$ or $s_{i}\subset\ell(c)$. It follows that
@@
is a finite disjoint union of components $z$ such that every $z$ is either a subinterval $a_{i},a_{j}]$ ($\subset[a,b]$) with $0\leq i<j\leq k$, where for non-closed $f$ the first and last inequality is strict, or a point in $(a,b)$. We call zzsimple component if the values $f(t)$ for $t$ immediately before and after $z$ lie on different sides of $\ell(c)$. Else, if they lie on the same side of $\ell(c)$, we call zzdouble component. We define
@@
We prove that the parity map $N$ is continuous on $S(f)$. Let $c\in S(f)$. We take a $\delta>0$ small enough such that the ball $B=B(c,\delta)\subset S(f)$ (Proposition 1.8) and that no vertical line that intersects $B$ contains a corner of $f[I]$ distinct from those whose preimages appear in $P(c)$. Let $c^{\prime}\in B$ with $c^{\prime}\neq c$. As we move $c$ to $c^{\prime}$, every simple component of $P(c)$ is preserved and every double component is either fixed or disappears or decays in two simple (point) components. It follows that $N(c)=N(c^{\prime})$. So $N$ is continuous at $c$ and on $S(f)$.
@@
Step 2. Two properties of the parity map $N$. The first property is the additivity of $N(c,f)$ in the variable $f$. Let $f_{1},f_{2}\colon I=[a,b]\to\mathbb{R}^{2}$ be two non-closed PL maps such that $f_{1}(b)=f_{2}(a)$. Let $f_{3}=f_{1}+f_{2}$ and let $c\in S(f_{1})\cap S(f_{2})\cap S(f_{3})$. Then
@@
Here is an outline of the proof of this equality: if $S_{i}$ is the set of simple components in $P(c,f_{i})$, $i=1,2,3$, then there is a bijection between
$S_{3}\,\text{ and }\,(S_{1}\times\{0\})\cup(S_{2}\times\{1\})\,.$
We leave details of the bijection to the reader as an exercise.
@@
To prove it, we consider the partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of the domain of $f$ and the components $z_{1}<z_{2}<\dots<z_{l}$ of
@@
As we know, every component $z_{i}$ is an interval $a_{r_{i}},a_{s_{i}}]$ ($\subset[a,b]$) with $0<r_{i}<s_{i}<k$ or a point $\{t_{i}\}$ with $t_{i}\in(a,b)$. It is clear that $P(c,f)\neq\emptyset$ and $l\geq 1$ (Proposition 1.22). We call the $l+1$ open subintervals
@@
of $(a,b)$ gaps. It follows from the assumption on $c$ that for every gap $g_{i}$ and every $t\in g_{i}$, all values $f(t)$ lie on only one side of $\ell(c)$, the left side $L$ or the right side $R$. We accordingly assign to $P(c,f)$ a word $w_{0}w_{1}\dots w_{l}\in\{L,R\}^{l+1}$. For $i=1,2,\dots,l$, the transition $w_{i-1}w_{i}$ with $w_{i-1}=w_{i}$ corresponds to a double component of $P(c,f)$, and that with $w_{i-1}\neq w_{i}$ to a simple component. Since $w_{0}\neq w_{l}$, the number of the latter transitions is odd.
@@
Step 3. Some two points in the complement of any circuit are separated. Let $f\colon I\to\mathbb{R}^{2}$ be a circuit and $D=\mathbb{R}^{2}\setminus f[I]$. We may assume that $f[I]$ lies to the right of the $y$-axis. Let $l\in f[I]$ and $p\in f[I]$ be, respectively, a leftmost and a rightmost point in $f[I]$ (Propositions 1.6 1.19). These points divide $f$ in two arcs $f_{1},f_{2}\colon I\to\mathbb{R}^{2}$ such that $f_{1}^{0}\cap f_{2}^{0}=\emptyset$, $f_{1}[I]\cup f_{2}[I]=f[I]$ and $\mathrm{ep}(f_{1})=\mathrm{ep}(f_{2})=\{l,p\}$ (Proposition 1.21). Clearly, $l_{x}<p_{x}$ (Proposition 1.20). Let $\ell$ be any vertical line with $x$-coordinate in $(l_{x},p_{x})$. Let $a\in f[I]\cap\ell$ be the highest point of this intersection (Proposition 1.19). We may assume that $a\in f_{1}^{0}$. Similarly, let $b\in f_{1}^{0}\cap\ell$ be the lowest point of this intersection. We take a point $c\in\ell$ below $b$ but such that
$0<d_{2}(c,\,b)<d_{2}(b,f_{2}[I])$
@@
— then $c\in D$. The point $c$ exists by Propositions 1.6 and 1.10, by which $d_{2}(b,f_{2}[I])>0$. We show that for every PL map $g\colon I\to\mathbb{R}^{2}$ that joins $c\in D$ to the origin $\overline{0}=\langle 0,0\rangle\in D$ ($f[I]$ lies to the right of the $y$-axis), $g[I]\cap f[I]\neq\emptyset$. In view of Propositions 1.8 and 1.18 it means that the complement $D=\mathbb{R}^{2}\setminus f[I]$ is disconnected.
@@
Suppose for the contrary that $g\colon I\to\mathbb{R}^{2}$ is a PL map joining $c$ to $\overline{0}$ in $D$, so that $g[I]\cap f[I]=\emptyset$. Let $\kappa$ ($\subset\mathbb{R}^{2}$) be the half-line going down from $c$ and let
$\lambda=c\,b\cup f_{3}[I]\cup\ell(a)\ \ (\subset\mathbb{R}^{2})\,$,
@@
where $cb$ is the straight segment joining these two points, $f_{3}$ is a subarc of $f_{1}$ joining $b$ to $a$ (if $b=a$, we omit $f_{3}[I]$) and, as before, $\ell(a)$ is the half-line going up from $a$. Let $h>0$ be the minimum of the three positive distances
$d_{2}(g[I],\,f[I])>0,\ d_{2}(\kappa,\,f_{1}[I])>0\,\text{ and }\,d_{2}(\lambda,\,f_{2}[I])>0\,.$
@@
The first distance is positive by the assumption that $g[I]\cap f[I]=\emptyset$ and by Propositions 1.6 and 1.10. The latter two distances are positive by Propositions 1.4, 1.5, 1.6, 1.7, and 1.11.
@@
We approximate the given circuit $f\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ sufficiently closely by a closed PL map. By Proposition 1.12, there exists a partition $\alpha=a_{0}<a_{1}<\dots<a_{k}=\beta$ of $\alpha,\beta]$ such that $\{a,b,l,p\}\subset\{f(a_{i})\colon\;i=0,1,\dots,k\}$ and
@@
We join for $i=1,2,\dots,k$ the consecutive points $f(a_{i-1})$ and $f(a_{i})$ ($f(a_{0})=f(a_{k})$) by straight segments and obtain a closed PL map $f_{4}\colon I=[\alpha,\beta]\to\mathbb{R}^{2}$ such that all corners of $f_{4}[I]$ lie on $f[I]$ and include the four points $a$, $b$, $l$, and $p$. Also, for every $z\in f_{4}[I]$ we have $d_{2}(z,f[I])<\frac{1}{2}h$. Let $f_{5}$, $f_{6}$, and $f_{7}$ be the non-closed PL sub-maps of $f_{4}$ spanned by the corners of $f_{4}[I]$ lying in $f_{1}[I]$, $f_{2}[I]$, and $f_{3}[I]$, respectively. The above definition of $h$ and the $\frac{1}{2}h$-approximation of $f_{4}[I]$ by $f[I]$ imply that
@@
The first equality follows from the fact that since $f[I]$ lies to the right of the $y$-axis, so does $f_{4}[I]$. The second equality follows from Step 1 because $f_{4}[I]\cap g[I]=\emptyset$. The third equality follows from the first property of $N$ in Step 2 because (we can define $f_{5}$ and $f_{6}$ in such a way that) $f_{4}=f_{5}+f_{6}$. We have $N(c,f_{5})=1$ by the second property of $N$ in Step 2 (since $f_{5}[I]\cap\kappa=\emptyset$, the point $c$ lies below all intersections of the line $x=c_{x}$ with $f_{5}[I]$). Finally, $N(c,f_{6})=N(a,f_{6})$ by Step 1 because $(cb\cup f_{7}[I])\cap f_{6}[I]=\emptyset$, and $N(a,f_{6})=0$ by the definition of $N$ because $\ell(a)\cap f_{6}[I]=\emptyset$.
@@ 4 Concluding remarks
Jordan’s proof of his theorem in [14] was often questioned in the past, and it was claimed that the first correct proof of the JT was given by Veblen [18]. Hales [8] reversed these assessments and showed that Jordan’s proof is basically correct when some gaps are filled, but that Veblen’s proof has serious problems. Filippov’s article [5] is preceded by the article [19] by A. I. Vol’pert, which contains a similar proof of the JT. The footnote on the first page of [5] reads “After this work was completed, the author learned that a proof of Jordan’s theorem, very close to the one in this article, was somewhat earlier found by A. I. Vol’pert (Lviv)”. Authors of these two proofs, Aleksei F. Filippov (1923–2006) and Aizik I. Volpert (Vol’pert) (1923–2006), were excellent mathematicians — see [2, 1].
In recent years, the ways in which we (mathematicians and philosophers) think about the links between informal and formal proofs, and about the role of formalized mathematics, have changed considerably and dramatically compared to the times before, say, 2000: see, for example, [3, 4, 6, 7]. What should be the results of the work of a research mathematician? A possible answer is: rigorous proofs of mathematical theorems. An informal proof is rigorous if it can be routinely translated into a formal proof ([11]), formal in the sense of the previously mentioned developments; see also [9, 10]. In this article, we wanted to lift the beautiful proofs of Thomassen and Filippov to the modern level of rigor, and we hopefully succeeded. The preprint [3] reports on the process of autoformalization of Munkres’ textbook [16] on topology. It is an impressive achievement; it seems that very soon AI will do everything for us. But does this autoformalization help a student interested in topology learn and understand ([12, 13]) the proofs of theorems in [16]? Not much. In contrast, we hope that our article will be helpful to anyone wishing to understand the two proofs of the WJT.
I lectured on Thomassen’s proof of the WJT in the course Mathematical Analysis 3 in the years 2023–2026. For the next year, I plan to switch to the proofs of the JT in [5, 19].